lim x + 2x2 − x3/ 3 − x2



1
Expert's answer
2022-04-06T15:44:20-0400
"\lim _{x \rightarrow \infty}\left(\frac{x+2 x^{2}-x^{3}}{3-x^{2}}\right)\\
\text{Divide by highest denominator power: } \frac{\frac{1}{x}+2-x}{\frac{3}{x^{2}}-1}\\[2mm]
\begin{aligned}
&=\lim _{x \rightarrow \infty}\left(\frac{\frac{1}{x}+2-x}{\frac{3}{x^{2}}-1}\right) \\
&\lim _{x \rightarrow a}\left[\frac{f(x)}{g(x)}\right]=\frac{\lim _{x \rightarrow a f} f(x)}{\lim _{x \rightarrow a g} g(x)}, \quad \lim _{x \rightarrow a g}(x) \neq 0
\end{aligned}
\\[2mm]
\textit{With the exception of indeterminate form}\\[2mm]
=\frac{\lim _{x \rightarrow \infty}\left(\frac{1}{x}+2-x\right)}{\lim _{x \rightarrow \infty}\left(\frac{3}{x^{2}}-1\right)}\\
\lim _{x \rightarrow \infty}\left(\frac{1}{x}+2-x\right)=-\infty\\
\lim _{x \rightarrow \infty}\left(\frac{3}{x^{2}}-1\right)=-1\\
=\frac{-\infty}{-1}\\
=\infty"